# Decentralized Identifier NSFW Not Safe For Work

### v1

@@ 1. Introduction
With the exclusion of the DIDs, any competing operator can build the above components, and users can freely select between them. For example, users can migrate their data between competing PDSes or configure the use of alternative Feed Generators. This plurality constitutes a radical shift from the walled-garden approach espoused by existing major players and proposes novel innovations with respect to other DSNs. Bluesky produces a complex system of interconnected components, operators, and users with a diversity of experiences. In this paper, we examine the operation of these components, study whether operators uptake the opportunity of providing them, whether this results in competing offerings and how users choose among offerings and operators. We particularly focus on how multiple operators manage the critical components: PDSes, Feed Generators, and Labelers. With this in mind, we gather the first large-scale Bluesky dataset, covering $5{,}523{,}919$ users, $225{,}461{,}969$ posts, $40{,}398$ Feed Generators, and $62$ Labelers.
@@ 2. Bluesky Primer
Bluesky aims to create a social network with a distinct set of characteristics. Each system component (e.g. data storage, moderation engine) is open-source and can be replicated, operated, and modified by the community. Furthermore, the network is decentralized. Users have control over their data, and can migrate between competing operators (e.g. PDS providers) freely. Lastly, Bluesky attempts to be scalable and easy to use, by avoiding redundant communication and providing a user experience similar to centralized social networks. In the rest of the paper, we use Bluesky to refer to the social network platform and Bluesky PBS when referring to the company developing the platform. In this section, we provide an overview of the key concepts of Bluesky.
@@
Each DID points to its associated DID Document (DID Document), a document that stores service information about the user. This includes the endpoint of the Personal Data Server (PDS) storing the user’s data repository, the user’s handle, as well as public keys to verify signatures on user content. There are currently two supported DID schemata: PLC and WEB. They differ in how the DID Document are retrieved: (1) for PLC DIDs, the associated document is downloaded from the plc.directory service, which is operated by Bluesky PBS. (2) WEB DIDs consist of a Fully-Qualified Domain Name (FQDN). The associated document must be located at https://<fqdn>/.well-known/did.json .
@@
Personal Data Servers. Repos are hosted by PDSes. A PDS can hold multiple repositories, but each repository is served by just one PDS that the user can choose. Bluesky PBS operates the default PDSes. However, it has recently become possible to self-host PDSes and federate with the network. [^1] https://docs.bsky.app/blog/self-host-federation Thus, users can migrate to different PDSes while maintaining their social graph. This requires updating the endpoint in the user’s DID Document. The PDSes also store user preferences, which are only accessible by the authenticated user.
The Relay. Checking for repo updates individually is resource-intensive, as each client would need to contact many different PDSes. To streamline the process, a centralized Relay aggregates user interactions across PDSes. This is a central store that replicates the repo data structures from all known PDSes. The relay then provides the Firehose — a real-time feed of all activity in the network that anyone can subscribe to. The Firehose has a retention time of three days and includes user repo updates as well as informational and service-related events (e.g. updates to user handles or DID Documents). Bluesky PBS runs the default Relay and Firehose at bsky.network. However, other providers could offer a competing service if they wish.
@@
Labelers operate as regular accounts, identified by a DID, with a repository for their activity. Each Labeler publishes a service information record in its repository, describing the values and default actions to be taken for its labels, e.g., to display a warning. Functionally, a labeling service implements an endpoint that publishes the labels. The endpoint is listed in the DID Document of the hosting account and is publicly accessible without authentication. Any entity can connect to this endpoint to retrieve the stream of labels produced. Bluesky PBS has been running the largest labeling service for the network. Recently, the ecosystem opened enabling anyone to run a Labeler. [^2] https://docs.bsky.app/docs/advanced-guides/moderation
@@
The AppView. The AppView collates the data produced across the network into a usable format, and makes it available to clients via a public API. In practice, the AppView consumes the Firehose and information from other system components (e.g. Labelers), stores them in a database, and provides the results in an easy way for a final app to display. There is currently one AppView for Bluesky, operated by Bluesky PBS.
@@ 3. Datasets
DID Documents and FQDN Handles. We then download the DID Documents (if available) for all user identifiers obtained in the previous step. Recall that these documents list fully qualified domain name (FQDN) handles, endpoints for users’ PDSes, and other service information. We obtain documents both from the centralized plc.directory service, [^3] https://plc.directory/ operated by Bluesky PBS, used as the default for account creations (the did:plc method), as well as .well-known/did.json paths via HTTPs (the did:web method) (see Section 2). We download the full snapshot over the course of one week in March 2024. This yields a total of 5,077,159 DID Documents from the PLC server, with an additional six using the did:web method. Next, we extract FQDN handles from the DID Documents for further analysis.
@@
Feed Post Dataset. We collect URIs of all Feed Generator posts from the Appview getFeed endpoint bi-weekly. We then correlate the URIs with the posts in the Repositories Dataset to obtain the full content of each post. Although 100% of Feed Generators with metadata were marked as both online and valid, we were only able to get posts for 93% of the Feed Generators. While there is a way to delete Feed Generator records from the repos, provided this record is still there, there is no way to distinguish between permanent and temporary unavailability. We thus exclude 4.3% of Feed Generators without metadata from our analysis.
@@ 4. User Activity
Growth. We observe the initial activity in users’ repositories dating back to November 2022, when Bluesky was launched (Figure 1). This coincides with Elon Musk’s acquisition of Twitter and the subsequent firing of half of the employees. The events following the takeover spurred interest in decentralized social media platforms (He et al., 2023). Bluesky experienced an large growth in daily user activity over the subsequent eight months following November 2022. The number of active users increased from mere hundreds in December 2022 to hundreds of thousands by July 2023. Another significant surge in daily user activity occurred in February 2024, when Bluesky transitioned from an invite-only platform to a public one. While Bluesky experienced initial growth, we also observe stagnation and even a decrease in daily active users. The number of daily active users decreased by $\approx$ 60K between March and May 2024.
@@
We further investigate the activity of $\approx$ 2M Bluesky users, who posted at least once, based on the self-assigned language tags attached to their posts (Figure 2). We verify a small ($\approx 0.1\%$) random portion of these posts and find that they indeed correspond to the language indicated in the tags. Language-specific communities roughly follow the global trend of user activity but we also notice some discrepancies. For instance, opening the platform to the public in February 2024 significantly increased the number of active Japanese-speaking users while the German-speaking community remained largely unaffected. In April, the Portuguese-speaking community experienced a sharp growth from $\approx$ 3K to $\approx$ 30K, most likely caused by direct marketing actions and recent documentation transalation (bsk, [n. d.]).
Current Status. As of April 2024, we observe a consistent presence of around 500K active users daily on the platform, contributing approximately 3M likes, 800K posts, and 300K reposts daily (Figure 1). The network size is comparable to Mastodon but significantly smaller than Twitter with 1M (mas, [n. d.]) and 245M (twi, [n. d.]) daily active users, respectively. While English remains the platform’s main language ($\approx$ 800K users), more than 700K users use Japanese language tags, suggesting a diverse user base. The platform has recognized the importance of the community, for instance, by implementing a dedicated kawaii mode. [^4] https://bsky.app?kawaii=true Additionally, Portuguese and German emerge as the next popular languages among users.
@@
Non-Bluesky content. We find that the Firehose also distributes content not covered by the Bluesky-related lexicons.These are records in user repositories targeted for a different AppView, i.e., Bluesky cannot decode or display them. We find $1{,}855$ events (out of $\approx 280$M) relating to these records. Among the targeted applications we found WhiteWind [^5] https://whtwnd.com/ to be the most popular public application. WhiteWind attempts to bring long-form blogging to ATProto. Users can log in using their Bluesky account and write articles using markdown. The articles are then saved in the user’s repo, hosted on their PDS. The WhiteWind AppView and Frontend can decode the entries and display them on the website.
These alternative applications require the Bluesky infrastructure (e.g. the Firehose) to index and re-publish non-Bluesky content. The ability to use already deployed infrastructure without additional cost is beneficial to the growth of the ecosystem. However, it remains to be seen whether this remains viable as those platforms grow.
@@
Validating Handle Ownership. Bluesky services currently recognize DIDs derived from either did:web, defined by the W3C Credentials Community Group (Gribneau et al., 2023), or did:plc, established by Bluesky PBS for Atproto (Holmgren et al., 2023). Interestingly, we identify only six did:web identities. This observation could stem from the accessibility of the mechanism supported by Bluesky PBS, along with the distinction from domain name handles, particularly in the immutability of the identifier. Unlike domain name handles, the did:web domains associated with a DID cannot be changed (Kleppmann et al., 2024).
@@
Registrar Concentration. Next, we perform a WHOIS scan to evaluate the concentration of registered domain names among various registrars identified using IANA IDs. We collect WHOIS data for 47,728 (92%) registered domain names and sucessfully extract the IANA ID for 39,403 (76%) domain names. We cannot retrieve the IANA ID for all registrars. While ICANN-accredited registrars must display the IANA IDs for new and legacy generic top-level domains (TLDs), public WHOIS records for country-code TLD (ccTLD) operators might not always contain this information. This could be because ccTLDs opt to provide limited data in WHOIS or because registrars are locally accredited by ccTLDs, exempting them from the requirement for ICANN-accredited registrars to have an IANA ID (Bayer et al., 2022).
We find a high degree of centralization of domain handles. While we identify 39,403 registered domains spread across 249 registrars, 50% of the domains are registered in just four registrars. Namecheap, an ICANN-accredited registrar, accounts for 21%. The collaboration announced in May 2023 between Bluesky and Namecheap to streamline the process of purchasing and linking domain names to Bluesky explains the latter. [^7] https://bsky.social/about/blog/7-05-2023-namecheap We argue that dedicated support from additional registrars will be necessary to avoid the risks of excessive registrars centralization in the future.
@@ 6. Content Moderation
Content moderation in Bluesky relies on two main elements: (1) Labelers, which produce short textual labels attached to objects, and (2) User preferences indicating (a) which Labelers to subscribe to (i.e. use labels from) and (b) how to react to a given label (e.g. by hiding the post). In this section, we focus on the Labelers producing the labels, and later inspect the labels themselves and the process of their issuance. Note, the user preferences are not publicly visible and we make no attempt to reveal them due to ethical considerations.
@@ 6.2. Labels
Looking at Table 4, the most-applied label for posts is no-alt-text, applied by the most popular community Labeler. This label marks posts with attached media missing an alternative text description for the media. The next most-frequently applied labels (porn and sexual) describe Not Safe For Work (NSFW) content and are almost exclusively applied by the official Bluesky Labeler. The tenor-gif label is applied by a community Labeler. It marks posts containing a reaction GIF from the popular Tenor [^10] https://tenor.com/ GIF keyboard.
@@
Another challenge is that there is no official list of potential label values to apply (apart from $7$ labels, of which some are exclusive to the official Bluesky Labeler as outlined earlier). As such, different Labelers use different labels with similar meanings. While this can pose issues regarding coherent labeling, it also offers flexibility to developers. Ultimately, Labelers play just one part in the overall moderation infrastructure: users choose which Labelers to subscribe to, and how their client should react to specific labels produced by these services. Because users choose which Labelers to subscribe to, there is a potential challenge in how to discover appropriate Labelers, and make informed decision about which labels to trust.

### v2

@@ Abstract.
Bluesky, Decentralized Social Networks, Social Network Analysis
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2024 ACM Internet Measurement Conference; November 4–6, 2024; Madrid, Spainbooktitle: Proceedings of the 2024 ACM Internet Measurement Conference (IMC ’24), November 4–6, 2024, Madrid, Spaindoi: 10.1145/3646547.3688407isbn: 979-8-4007-0592-2/24/11ccs: Networks Network measurementccs: Networks Social media networksccs: Information systems Social networks
@@ 1. Introduction
Any competing operator can build the above components, and users can freely select between them. For example, users can migrate their data between competing PDSes or configure the use of alternative Feed Generators. This plurality constitutes a radical shift from the walled-garden approach espoused by existing major players and proposes novel innovations with respect to other DSNs. Bluesky produces a complex system of interconnected components, operators, and users with a diversity of experiences. In this paper, we examine the operation of these components, study whether operators uptake the opportunity of providing them, whether this results in competing offerings and how users choose among offerings and operators. We particularly focus on how multiple operators manage the critical components: PDSes, Feed Generators, and Labelers. With this in mind, we gather the first large-scale Bluesky dataset, covering $5{,}523{,}919$ users, $225{,}461{,}969$ posts, $40{,}398$ Feed Generators, and $62$ Labelers.
@@ 2. Bluesky Primer
Bluesky is a decentralized network where each system component (e.g. data storage, moderation engine) is open-source and can be replicated, operated, and modified by the community. Users have control over their data, and can migrate between competing operators freely. Bluesky attempts to be scalable and easy to use, by avoiding redundant communication and providing a user experience similar to centralized social networks. We use Bluesky to refer to the social network platform and Bluesky PBC when referring to the company developing the platform. In this section, we provide an overview of the key concepts of Bluesky.
@@
Each DID points to its associated DID Document (DID Document), a document that stores service information about the user. This includes the endpoint of the Personal Data Server (PDS) storing the user’s data repository, the user’s handle, as well as public keys to verify signatures on user content. There are currently two supported DID schemata: PLC and WEB. They differ in how the DID Document are retrieved: (1) for PLC DIDs, the associated document is downloaded from the plc.directory service, which is operated by Bluesky PBC. (2) WEB DIDs consist of a Fully-Qualified Domain Name (FQDN). The associated document must be located at https://<fqdn>/.well-known/did.json .
@@
Personal Data Servers. Repos are hosted by PDSes. A PDS can hold multiple repositories, but each repository is served by just one PDS that the user can choose. Bluesky PBC operates the default PDSes. However, it has recently become possible to self-host PDSes and federate with the network. [^1] https://docs.bsky.app/blog/self-host-federation Thus, users can migrate to different PDSes while maintaining their social graph. This requires updating the endpoint in the user’s DID Document. The PDSes also store user preferences, which are only accessible by the authenticated user.
The Relay. Checking for repo updates individually is resource-intensive, as each client would need to contact many different PDSes. To streamline the process, a centralized Relay aggregates user interactions across PDSes. This is a central store that replicates the repo data structures from all known PDSes. The relay then provides the Firehose — a real-time feed of all activity in the network that anyone can subscribe to. The Firehose has a retention time of three days and includes user repo updates as well as informational and service-related events (e.g. updates to user handles). Bluesky PBC runs the default Relay and Firehose at bsky.network. However, other providers could offer a competing service if they wish.
@@
Labelers operate as regular accounts with a repository for their activity. Each Labeler publishes a service information record in its repository, describing the values and default actions to be taken for its labels, e.g. to display a warning. Functionally, a labeling service implements an endpoint that publishes the labels. The endpoint is listed in the DID Document of the hosting account and is publicly accessible without authentication. Any entity can connect to this endpoint to retrieve the stream of labels produced. Recently, the ecosystem opened enabling anyone to run a Labeler. [^2] https://docs.bsky.app/docs/advanced-guides/moderation
@@
The AppView. The AppView collates the data produced across the network into a usable format, and makes it available to clients via a public API. In practice, the AppView consumes the Firehose and information from other system components, stores them in a database, and provides the results in an easy way for a final app to display. There is currently one Bluesky AppView, operated by Bluesky PBC.
@@ 3. Datasets
DID Documents and FQDN Handles. We then download the DID Documents (if available) for all user identifiers obtained in the previous step. Recall that these documents list fully qualified domain name (FQDN) handles, endpoints for users’ PDSes, and other service information. We obtain documents both from the centralized plc.directory service, [^3] https://plc.directory/ operated by Bluesky PBC, used as the default for account creations (the did:plc method), as well as .well-known/did.json paths via HTTPs (the did:web method) (see Section 2). We download the full snapshot over the course of one week in March 2024. This yields a total of 5,077,159 DID Documents from the PLC server, with an additional six using the did:web method. We extract FQDN handles from the DID Documents for further analysis.
@@
Feed Post Dataset. We collect URIs of all Feed Generator posts from the Appview getFeed endpoint bi-weekly. We then correlate the URIs with the posts in the Repositories Dataset to obtain the full content of each post. Although 100% of Feed Generators with metadata were marked as both online and valid, we were only able to get posts for 93% of them. While there is a way to delete Feed Generator records from the repos, provided this record is still there, there is no way to distinguish between permanent and temporary unavailability. We thus exclude 4.3% of Feed Generators without metadata from our analysis.
@@ 4. User Activity
First, we analyze the user activity on Bluesky. We study the growth dynamics of the platform, the most popular accounts, and its current size compared to other, established social networks. This allows us to understand the platform’s current status and its user base. We deliberately omit a deeper analysis of the social graph, as Bluesky does not introduce novel solutions in this regard. Finally, we investigate whether the theoretically flexible ATProto infrastructure is used to distribute content not related to the social network itself.
@@
Growth. We observe the initial activity in users’ repositories dating back to November 2022, when Bluesky was launched (Figure 1). This coincides with Elon Musk’s acquisition of Twitter and the subsequent firing of half of the employees. The events following the takeover spurred interest in decentralized social media platforms (He et al., 2023). Bluesky experienced a large growth in daily user activity over the subsequent eight months following November 2022. The number of active users increased from mere hundreds in December 2022 to hundreds of thousands by July 2023. Another significant surge in daily user activity occurred in February 2024, when Bluesky transitioned from an invite-only platform to a public one. While Bluesky experienced initial growth, we also observe stagnation and even a decrease in daily active users. The number of daily active users decreased by $\approx$ 60K between March and May 2024.
@@
We further investigate the activity of $\approx$ 2M Bluesky users, who posted at least once, based on the self-assigned language tags attached to their posts (Figure 2). We verify a small ($\approx 0.1\%$) random portion of these posts and find that they indeed correspond to the language indicated in the tags. Language-specific communities roughly follow the global trend of user activity but we also notice some discrepancies. For instance, opening the platform to the public in February 2024 significantly increased the number of active Japanese-speaking users while the German-speaking community remained largely unaffected. In April, the Portuguese-speaking community experienced a sharp growth from $\approx$ 3K to $\approx$ 30K, most likely caused by direct marketing actions and recent documentation translation (bsk, [n. d.]).
Current Status. As of April 2024, we observe a consistent presence of around 500K active users daily on the platform, contributing approximately 3M likes, 800K posts, and 300K reposts daily (Figure 1). The network size is comparable to Mastodon but significantly smaller than Twitter with 1M (mas, [n. d.]) and 245M (twi, [n. d.]) daily active users, respectively. While English remains the platform’s main language ($\approx$ 800K users), more than 700K users use Japanese language tags, suggesting a diverse user base. The platform has recognized the importance of the community, for instance, by implementing a dedicated kawaii mode. [^4] https://bsky.app?kawaii=true Additionally, Portuguese and German emerged as the next popular languages among users.
@@
Non-Bluesky content. We find that the Firehose also distributes content not covered by the Bluesky-related lexicons.These are records in user repositories targeted for a different AppView, i.e., Bluesky cannot decode or display them. We find $1{,}855$ events (out of $\approx 280$M) relating to these records. Among the targeted applications, we find WhiteWind [^5] https://whtwnd.com/ to be the most popular public application. WhiteWind attempts to bring long-form blogging to ATProto. Users can log in using their Bluesky account and write articles using markdown. The articles are then saved in the user’s repo, hosted on their PDS. The WhiteWind AppView and Frontend can decode the entries and display them on the website.
These alternative applications require the Bluesky infrastructure (e.g. the Firehose) to index and re-publish non-Bluesky content. The ability to use already deployed infrastructure without additional cost is beneficial to the growth of the ecosystem. Due to the current limited size of those applications, they have a limited impact on the infrastructure. However, it remains to be seen whether this remains viable as those platforms grow.
Takeaways. Bluesky has experienced significant growth since its launch in November 2022. However, the growth has been driven mostly by specific events such as the public launch and stagnates between those events. Bluesky is still significantly smaller than its centralized counterpart and does not yet show signs of the snowball effect or mass user migration. This is further confirmed by the current lack of popular celebrity or institutional accounts. At the same time, the good user experience and a rapidly increasing number of features could enable Bluesky to attract specific country- or interest-specific communities. This is exemplified by a recent mass migration of Brazilian users after X/Twitter was banned in the country (CNBC, [n. d.]). [^6] Unfortunately, the migration happened outside of our data collection period. Finally, opening the platform infrastructure to third-party applications could further boost the growth and diversity of the ecosystem but for now, remain in their infancy.
@@
Validating Handle Ownership. Bluesky services currently recognize DIDs derived from either did:web, defined by the W3C Credentials Community Group (Gribneau et al., 2023), or did:plc, established by Bluesky PBC for Atproto (Holmgren et al., 2023). Interestingly, we identify only six did:web identities. This observation could stem from the accessibility of the mechanism supported by Bluesky PBC, along with the distinction from domain name handles, particularly in the immutability of the identifier. Unlike domain name handles, the did:web domains associated with a DID cannot be changed (Kleppmann et al., 2024).
@@
Registrar Concentration. Next, we perform a WHOIS scan to evaluate the concentration of registered domain names among various registrars identified using IANA IDs. We collect WHOIS data for 47,728 (92%) registered domain names and successfully extract the IANA ID for 39,403 (76%) domain names. We cannot retrieve the IANA ID for all registrars. While ICANN-accredited registrars must display the IANA IDs for new and legacy generic top-level domains (TLDs), public WHOIS records for country-code TLD (ccTLD) operators might not always contain this information. This could be because ccTLDs opt to provide limited data in WHOIS or because registrars are locally accredited by ccTLDs, exempting them from the requirement for ICANN-accredited registrars to have an IANA ID (Bayer et al., 2022).
We find a high degree of centralization of domain handles. While we identify 39,403 registered domains spread across 249 registrars, 50% of the domains are registered in just four registrars. Namecheap, an ICANN-accredited registrar, accounts for 21%. The collaboration announced in May 2023 between Bluesky and Namecheap to streamline the process of purchasing and linking domain names to Bluesky explains the latter. [^8] https://bsky.social/about/blog/7-05-2023-namecheap We argue that dedicated support from additional registrars will be necessary to avoid the risks of excessive registrar centralization in the future.
@@
Takeaways. Bluesky offers a variety of options for identity management ranging from custodial subdomain handles to self-managed domain names. Unsurprisingly, the vast majority of users opt for the convenience of automatic subdomain handles under bsky.social that resembles traditional social media authentication. This leaves Bluesky in full control over user identities but also enables the platform to avoid the trap of fully decentralized, complex identity management that could deter regular users. At the same time, institutions or tech-savvy users have the option to manage their keys and link their accounts to their domain names making their identity largely independent of the Bluesky operators. We observe that more decentralized options to identity management can attract users when made more accessible. It remains to be seen whether such a flexible approach to identity management will enable the onboarding of a large number of users with the centralized approach while later allowing them to take more control over their identity as the decentralized identity solutions mature and develop user-friendly interfaces.
@@ 6. Content Moderation
Content moderation is a crucial aspect of any social media platform. The traditional social media platforms have faced criticism for their opaque and inconsistent moderation practices. Bluesky presents a novel approach that relies on two main elements: (1) Labelers, which produce short textual labels attached to objects, and (2) User preferences indicating (a) which Labelers to subscribe to (i.e. use labels from) and (b) how to react to a given label (e.g. by hiding the post). In this section, we focus on the Labelers producing the labels, and later inspect the labels themselves and the process of their issuance. We try to reveal how the moderation process is split between the official Bluesky Labeler and community-run Labelers, and whether there is a significant overlap between them. Note, the user preferences are not publicly visible and we make no attempt to reveal them due to ethical considerations.
@@ 6.2. Labels
Looking at Table 4, the most-applied label for posts is no-alt-text, applied by the most popular community Labeler. This label marks posts with attached media missing an alternative text description for the media. The next most frequently applied labels (porn and sexual) describe Not Safe For Work (NSFW) content and are almost exclusively applied by the official Bluesky Labeler. The tenor-gif label is applied by a community Labeler. It marks posts containing a reaction GIF from the popular Tenor [^11] https://tenor.com/ GIF keyboard.
@@
Another challenge is that there is no official list of potential label values to apply (apart from $7$ labels, of which some are exclusive to the official Bluesky Labeler as outlined earlier). As such, different Labelers use different labels with similar meanings. While this can pose issues regarding coherent labeling, it also offers flexibility to developers. Note that Labelers have to provide descriptive metadata when declaring the list of label values they emit. Ultimately, Labelers play just one part in the overall moderation infrastructure: users choose which Labelers to subscribe to, and how their client should react to specific labels produced by these services. Because users choose which Labelers to subscribe to, there is a potential challenge in how to discover appropriate Labelers, and make informed decisions about which labels to trust.
@@ 6.3. Label issuance
Takeaways. Bluesky combines mandatory [^12] We note that, theoretically, one could run an alternative AppView component that would ignore labels issued by Bluesky PBC. platform-run moderation with a plethora of custom, user-led Labelers.The open Labeler ecosystem is still in its infancy and mostly issues specific labels (e.g. GIFs from a specific platform) avoiding more controversial topics (e.g. fake news or hate speech). Notably, the ease of running a Labeler and the flexibility of the system already enable it to be used not only for specific, niche topics, but also for downstream content recommendation (cf. Section 7). While we observe a high degree of automation in the label issuance process, some labels are still likely applied manually. This is especially true for the more subtle labels, such as threat or intolerant. We expect more labels to be issued automatically in the future as the content rate increases and technology matures.
As Bluesky grows, the platform might be also exposed to increasing regulatory pressure to moderate content.131313cf. https://bsky.app/profile/aaron.bsky.team/post/3l3gerugkbt27 for a recent example of increased need for moderation. This creates the risk of running forced, centralized moderation before the content is distributed to the network and limiting the role of decentralized Labelers. We note that it is already possible for Bluesky PBC to perform “infrastructure takedowns”, instantly removing data from their infrastructure in cases of clearly illegal content. While running a Labeler is relatively easy, the incentives for their operators to scale up their operations are unclear. In the long run, the decentralized Labelers might thus cover the issuance of high-quality labels for niche use cases, while the centralized moderation would handle the bulk of the sensitive content (e.g. CSAM) and remove it from the network.